IE 4355 Facilities Planning Homework Assignment - Capacity Planning Spring 2019 1. Under normal conditions, an LED flat panel manufacturer has the capacity of 60 panels/hour over its 40 hour work week. During periods of high demand it is able to allocate overtime to its work force so as to gain the equivalent of an extra 10 hours of production each week. A. What is the manufacturers effective capacity? B. What is the manufacturers design capacity? C. If the manufacturer had component supply problems resulting in an output of 55 panels/hour during a period of normal demand, what was their effective utilization during this time? D. If the manufacturer had an output of 73 panels/hour during a period of high demand (overtime was used), what was their effective utilization during this time?
Answer #1

A

In case of high demand the manufacturer can produce 60 panels/hour. However the number of hours per week becomes 50 hours. This means the effective capacity is 60*50 = 3000 units per week.

B

The design capacity is the normal capacity that was intended. For a 40 hour week, the capacity is 40*60 = 2400 units per week.

C

Utilization is given by actual output divided by design capacity. In this case the actual output is 55*40 = 2200. This means the effective utilization is 2200/2400 = 0.9166 or 91.66%

D

In this case, we know that during high level of demand, we can product 3000 units per week. However the output here is 73*50 = 3650 units. This makes the effective utilization as 3650/3000 = 1.2166 or 121.66%